Don’t Cry Over Spilled​ Milk, Inc., a manufacturer of​ mops, uses the weighted average method in its processing costing system.​Don’t Cry uses a traditional costing system in the allocation of its manufacturing overhead​ (MOH) to​ production; MOH costs are allocated to production on the basis of machine hours logged during the period. Total estimated MOH for the period was ​$400,000 and total estimated machine hours were​ 100,000. The following data concern the operations of the​ company’s first processing​department, Assembly, for June 2018.

Beginning Work in Process 800 mops, 60% complete with respect to direct materials, 10% complete with respect to conversion
Costs in Beginning Work in Process $3712 (of this amount, $1296 is for direct materials)
Units Started in June $16,000 mops
Direct Product Costs Added During June direct materials $47,076; direct labor $216,213
Ending Work in Process 300 mops, 60% completed with respect to direct materials, 70% completed with respect to conversion

During June, 78,648 actual machine hours were logged during production.

(round all calculation to 2 decimal places. all answers should include commas and 2 decimal places i.e.​ 1,000.00 not

1000.​)

1) The cost to assemble one mop in June was $____.

2) The cost of goods completed and transferred out of the Assembly department during June equals $_____.

3) The conversion costs included in Work in​ Process-Assembly at the beginning of July would equal $______.
Answer #1

Predetermined Overhead Rate = Total Estimated Manufacturing Overhead / Total Estimated Machine Hours

Total Estimated Manufacturing Overhead = $ 400,000

Total Estimated Machine Hours = 100,000 Hours

Predetermined Overhead Rate = 400,000 / 100,000

Predetermined Overhead Rate = $ 4 per Machine Hour

Overhead applied during June = Actual Machine Hour during June * Predetermined Overhead Rate

Overhead applied to Production during June = 78,648 * 4 = $ 314,592

Units Transferred during the Year = Opening Work in Process + Units Introduced during the Year - Ending Work in Process

Units Transferred Out during the Year = 800 + 16,000 - 300 = 16,500 Mops

Particulars Units % Materials EUP Materials % Conversion EUP Conversion
Units Transferred Out during the Year 16,500 100% 16,500 100% 16,500
Ending Work in Process Inventory 300 60% 180 70% 210
Total 16,800 16,680 16,710

Notes :-

EUP Stands for Equivalent Units of Production.

EUP Materials = Units * % Materials

EUP Conversion = Units * % Conversion

Total Material Costs = Beginning Work in Process Material Cost + Material Added during the Year

Total Material Costs = 1,296 + 47,076 = $ 48,372

Total Conversion Costs = Beginning Work in Process Conversion Costs + Direct Labour Cost + Overhead applied

Total Conversion Costs = 2,416 + 216,213 + 314,592 = 533,221

Cost per Equivalent Unit of Material = Total Material Costs / Equivalent Units for Materials

Cost per Equivalent Unit of Material = 48,372 / 16,680 = $ 2.90

Cost per Equivalent Unit of Conversion Costs = Total Conversion Costs / Equivalent Units for Conversion Costs

Cost per Equivalent Unit of Conversion Costs = 533,221 / 16,710 =$ 31.91

Question 1

Cost to assemble One Mop = Cost per Equivalent Unit for Materials + Cost per Equivalent Unit for Conversion Costs

Cost to assemble One Mop = 2.90 + 31.91 = $ 34.81

Question 2

Cost of Units Transferred Out during June = 16,500 Units * $ 34.81 per Unit

Cost of Unity Transferred Out during the June = $ 574,365

Question 3

Cost of Beginning Units for the Month of July = Cost of Beginning Inventory + Material Cost Added during the Year + Conversion Costs Added during the Year - Cost of Units Transferred Out during the Year

Cost of Beginning Unit for the Month of July = 3,712 + 47,076 + 216,213 + 314,592 - 574,369

Cost of Conversion Costs included Work in Process for the Month of July =$ 7,224

Ending WIP

Material = 180 Units * $ 2.90 = $ 522

Conversion Costs = 210 Units * $ 31.91 = $ 6,702